OpenStudy (anonymous):

Using the formula x = 2a + y, solve for y

OpenStudy (anonymous):

subtract 2a from both sides

OpenStudy (anonymous):

to isolate y by itself on the right side of the equation

OpenStudy (anonymous):

x - 2a = 2a + y - 2a x - 2a = y y = x - 2a

OpenStudy (anonymous):

y = x-2a because you move the variable you want to become isolated.